Sunday, January 9, 2011

Nightlog 2011-01-08

SA: Paul
SO: Siphelo
Others: Garith, Ardhisha, Charl, Peter


- The seeing was not good enough for IQ observations. We had cloud around at
the start of the evening, which returned just after 2.

- The tracker A-G motor was overheated when we got to the telescope. Charl and
Ardhisha came up to look at the alignment issue we had last night and also took
care of the overheating. The tracker behaved for the rest of the night.

- We had some alignment issues with the reference mirror not being found. Peter
came up to look at it.


No comments:

Post a Comment